# Emerging Activity Temporal Hypergraph (EATH), a model for generating realistic time-varying hypergraphs

### v1

@@ IV.2 Higher-order temporal and topological properties
Moreover, the EATH model captures the empirical structure both for the higher-order and pairwise topology at the aggregated level. The generated surrogate hypergraph replicates indeed well the empirical distributions of the hyperdegree and hyperstrength in the aggregated hypergraph $\mathcal{H}$, $P(D)$ and $P(S)$, both qualitatively and quantitatively (Fig. 5c,d). This also holds when considering the corresponding distributions in the projected graph $\mathcal{G}$ (Fig. 5e,f) and the hyperdegree distribution for specific group sizes (see SM). The EATH model also reproduces the evolution of the distributions of the interactions weights and of their heterogeneity (see SM), obtained considering the hypergraph aggregated from time 00 to time $t$. We note that the topology generated without the memory effects (EATHw) is very different: the average hyperdegree and projected degree are much higher, the distributions span a much broader interval, and the (hyper)degree reach in particular very large values; some nodes participate in a very large number of groups, having a large neighborhood, and only few groups and links are recurring (see Fig. 5 and SM). The empirical higher-order structure is thus reproduced by the combination of the memory mechanisms and the heterogeneity in node activity, as also noted in [12]. Their combined effect also allows to reproduce modularity values of the pairwise projection of the aggregated hypergraph close to the ones observed in the data, while the EATHw model leads always to very small modularity values (see SM).
@@
Figure 7: Generated nodes and hyperedges heterogeneity. Panels a and b show respectively the distributions of the instantaneous activity, $P(a_{h}(i))$, and of the persistence activity, $P(a_{T}(i))$, extracted from the empirical dataset (which are provided as input to the generation) and measured in the generated hypergraphs. In panel c (resp. f) each point corresponds to a node (resp. hyperedge) and we show the scatterplot between the burstiness of their inter-event time distributions, $\Delta B_{\tau}$, and of their event duration distributions, $\Delta B_{T}$. The gray dashed lines are the reference $\Delta B=0$. Panels d, e (resp. g, h) show the distributions $P(\Delta B_{T})$ and $P(\Delta B_{\tau})$ for nodes (resp. hyperedges): we consider only nodes and hyperedges with at least $10$ events. Note that in panels g and h there are no results for the EATHw case, since no group activated enough times to define its burstiness parameter. In all panels we consider the SFHH dataset and we compare it with hypergraphs generated using the model with (EATH) and without memory (EATHw), with the parameters extracted from the empirical hypergraph.
@@
Figure 6 focuses on the distributions of the quantities characterizing the dynamics of nodes and groups, at the global level: the dynamics generated for groups and nodes is heterogeneous, bursty and correlated, as empirically observed (see also the SM for the burstiness $\Delta B$ of the generated distributions). We note that, at the node level, the distributions of the number of events in trains of events and the inter-event times are well replicated even with the memory-less model EATHw. Indeed, the interaction dynamics of nodes is largely governed by their activity dynamics, which is built in the first step of the model (without reference to memory mechanisms). However, the EATHw generation process produces temporal dynamics that do not reproduce the empirical distributions of event durations for nodes nor any of the distributions characterizing the dynamics of groups (see Fig. 6). In this case indeed, the active groups are selected at each time-step only according to the activity of the nodes, thus it is not possible to obtain long durations for group interactions nor temporal correlations. Memory mechanisms are thus responsible for the generation of heterogeneous group durations, and allow also to reactivate the same groups at short intervals, creating trains of events [38, 43].
@@
Figure 8: Generated nodes participation to different group sizes. We consider the empirical SFHH hypergraph (Data) and the hypergraphs generated by the model with (EATH) and without memory (EATHw), with parameters extracted from the empirical hypergraph. In each case, we rank the nodes at each order $m$ according to the time spent interacting at that order. We then show the Pearson’s correlation coefficient $\rho(m,m^{\prime})$ between the rankings at order $m$ and $m^{\prime}$ (a,c,e). We show the evolution of the fraction $n_{f}^{m}(t)$ of active nodes among the ones occupying the top $fN$ position of the ranking at order $m$, for different $m$ (see legend), and $f=0.1$, focusing on the same time window (b,d,f). We also plot the total fraction $N_{t}/N$ of active nodes in the population. Panel g shows the distributions of the node participation ratios $P(y)$ (the red dashed line represents $y=1/(M-1)$).
@@
Finally, we highlight a limitation of the model. Not surprisingly given the creation rules described above, the surrogate temporal hypergraphs generated do not reproduce the dynamics of group aggregation and disaggregation (see SM): indeed the model’s group selection mechanisms only consider processes in which a group can lose or gain a single node. However, the empirical dynamics is more complex, with merging and splitting of groups of different sizes, as described in [11]. To reproduce such dynamics, additional mechanisms should be introduced in the model, with rules making it possible for a group at time $t-1$ to lose or gain several nodes at once.
@@ IV.3 Higher-order SIR dynamics
[…]
The goal of the EATH model is to produce surrogate temporal hypergraphs that can be used to simulate dynamical processes on hypergraphs. We therefore consider here one such process, namely a prototypical spreading process including higher-order mechanisms, and compare its dynamics and outcome when simulated on either the empirical or the surrogate hypergraphs. Specifically, we consider the higher-order SIR (susceptible-infected-removed) epidemic process based on a non-linear infection rate [47], which assumes a higher probability of contagion in case of simultaneous exposure to multiple sources of infection [47, 16]. Each node of the hypergraph can be susceptible, $S$, infected, $I$, or recovered, $R$. The dynamics evolves in discrete time, with a time-step $\delta t$ equal to the resolution of the hypergraph =$\mathscr{H}=\{(\mathcal{V}_{t},\mathcal{E}_{t})\}_{t=1}^{T/\delta t}$. At each time-step $t$, a susceptible node $j\in\mathcal{V}_{t}$ that is part of a hyperedge $e\in\mathcal{E}_{t}$ with $i_{e,t}$ infected nodes, gets infected with probability $(1-e^{-\lambda i_{e,t}^{\nu}})$ within the group $e$, where $\lambda\in[0,1]$ and $\nu\geq 1$. Hence, overall a susceptible node $j$ at time $t$ is infected with probability:
@@
Figure 9 shows the results of the simulations performed on both empirical and surrogate hypergraphs for the SFHH dataset (see SM for other datasets). The SIR dynamics on the EATH model replicates well the empirical one for all the parameters configurations, both in terms of the total impact of the epidemic $R_{\infty}$ and of the initial spreading $R_{0}$ (see Fig. 9a-f). Furthermore, the epidemic evolution on the EATH model is similar to the empirical one, replicating the timing and incidence levels of the epidemic peaks, despite a slight overestimation of the peak heights (see Fig. 9g,h). Simulations on the hypergraphs generated without memory mechanisms (EATHw) yield a much faster dynamics in the early stages, with a much higher first peak and overall a substantial overestimation of the impact of the spread.
@@ V EATH artificial and hybrid hypergraphs
Figure 10: Hybrid hypergraphs - mixing WS16 (system) and Utah_elem (population). We consider two hybrid hypergraphs (Hybrid_sub and Hybrid_uni) generated with the EATH model considering the properties of the datasets WS16 (conference context; $d_{1}$ system properties) and Utah_elem (school context; $d_{2}$ population properties). a-f: distributions of inter-event times, durations and numbers of events in a train of events, for nodes (a-c) and hyperedges (d-f); g,h: normalized hyperdegree distribution in the aggregated hypergraph, $P(D/\langle D\rangle)$, and normalized degree distribution in the projected aggregated graph, $P(D_{proj}/\langle D_{proj}\rangle)$; i-l: Pearson’s correlation coefficient $\rho(m,m^{\prime})$ between the node rankings at size $m$ and $m^{\prime}$, obtained considering the time spent by nodes interacting at each order; m-r: epidemic final-size $R_{\infty}$ and basic reproduction number $R_{0}$ as a function of the SIR epidemiological parameters $(\lambda,\mu)$, obtained in the same conditions of Fig. 9; s-t: scatterplots of the burstiness of inter-event time distributions $\Delta B_{\tau}$ and that of duration distributions $\Delta B_{T}$ for single nodes and hyperedges, considering only the nodes and hyperedges with at least 10 events.
The interest and use of the EATH model actually go beyond the generation of surrogate temporal hypergraphs. The model parameters and mechanisms governing the different generative phases can be tuned in order to generate hypergraphs with various properties. This opens the possibility to build temporal hypergraphs with arbitrary features, possibly even different from the empirical ones, and to test their impact, e.g. on dynamical processes. At the context/system level, we can for instance build a hypergraph: (i) with a different temporal activity dynamics, imposing an arbitrary modulation $\Lambda_{t}$; (ii) with groups of different sizes, by changing the distribution $\Psi(m)$; (iii) with different temporal and interaction scales, e.g. a longer time interval $T$ or a higher connectivity $\langle E_{t}\rangle$. At the population level, we can moreover: (i) consider a larger population $N$; (ii) consider arbitrary distributions of the activities, $a_{h}(i)$ and $a_{T}(i)$; (iii) set different order propensities $\varphi_{i}(m)$. In the SM we generate such cases and explore how these modifications affect the temporal, topological and dynamical features of the system.
@@
(i) Hybrid substitution (Hybrid_sub). The hypergraph is generated with the system parameters of $d_{1}$ and the population parameters of $d_{2}$. This case models a population of a specific environment (e.g. students of a school), which preserves its features but interacts in a different context (e.g. a conference), with different temporal patterns, activity levels and sizes of interactions. The connection between the system and population properties requires modifying the order propensity $\varphi_{i}(m)$ according to the maximum interactions sizes: if $M_{2}\geq M_{1}$, $\varphi_{i}(m)$ is kept as in $d_{2}$; if $M_{2}<M_{1}$ we first fix $\varphi_{i}(m)=\varphi_{i}(M_{2})\forall m\in[M_{2}+1,M_{1}]$, so that nodes preserve their propensity for interactions at large sizes and we then normalize the resulting $\varphi_{i}(m)$ on the interval $[2,M_{1}]$. We keep the level of interactions $\langle E_{t}\rangle=\langle E_{t,1}\rangle$, even if $N_{2}\neq N_{1}$, to focus on the pure effect of a population that interacts in a different environment, possibly characterized by a different interaction level;
@@
In Fig. 10 we show the properties of hybrid hypergraphs generated by mixing a conference (WS16, $d_{1}$) and a school dataset (Utah_elem, $d_{2}$). The generated hypergraphs differ from the empirical ones, but they still present realistic statistical properties for all the measures considered: a heterogeneous and bursty dynamics with temporal correlations both at the node and group level (Fig. 10 a-f), a realistic shape of the hyperdegree and degree distributions in the aggregated hypergraph and in its pairwise projection (Fig. 10 g,h), heterogeneity in the behaviours of the nodes both in the propensity to interact at different orders (Fig. 10 i-l) and temporally (Fig. 10 s,t). The outcome of a SIR dynamics presents as well a realistic pattern even if some differences are observed with respect to the original dataset $d_{1}$ (Fig. 10 m-r): in Hybrid_sub the epidemic is less effective, since the average number of interactions per node is lower; in Hybrid_uni the dynamics is more similar, but still differs due to the presence of $d_{2}$ nodes with different behaviours. In the SM we also consider other hybrid hypergraphs.
@@ VI Discussion
The EATH model can generate surrogate temporal hypergraphs emulating empirical systems and their different levels of heterogeneity. It captures and reproduces a wide set of empirical temporal and topological properties. Moreover, the EATH model is able to replicate the higher-order properties of the system that shape higher-order infection dynamics, generating thus surrogate temporal hypergraphs useful to investigate contagion processes.
@@ VII.1 Datasets and pre-processing
We consider several publicly available datasets of face-to-face interactions collected by different collaborations. The datasets consist in pairwise interactions collected with a resolution of 20 seconds through RFID proximity sensors. We consider six datasets collected by the SocioPatterns collaboration [49, 50, 51] describing interactions: in a hospital (LH10 [53], with $N=75$ and $T=96$ hours), in an elementary school (LyonSchool [54], with $N=242$ and $T=2$ days), in a high-school (Thiers13 [55], with $N=327$ and $T=5$ days), in three conferences (SFHH [50], with $N=403$ and $T=2$ days; ICCSS17 [56], with $N=274$ and $T=4$ days; WS16 [56], with $N=138$ and $T=2$ days); and two datasets collected by the Contacts among Utah’s School-age Population (CUSP) project [57], describing interactions in an elementary school (Utah_elem [57], with $N=339$ and $T=2$ days) and in a middle school (Utah_mid [57], with $N=591$ and $T=2$ days). These datasets describe social interactions in different environments, mediated by different mechanisms and activities.

### v2

@@ IV.2 Higher-order temporal and topological properties
Moreover, the EATH model captures the empirical structure both for the higher-order and pairwise topology at the aggregated level. The generated surrogate hypergraph replicates indeed well the empirical distributions of the hyperdegree and hyperstrength in the aggregated hypergraph $\mathcal{H}$, $P(D)$ and $P(S)$, both qualitatively and quantitatively (Fig. 5c,d). This also holds when considering the corresponding distributions in the projected graph $\mathcal{G}$ (Fig. 5e,f) and the hyperdegree distribution for specific group sizes (see SM, Sec. IB). The EATH model also reproduces the evolution of the distributions of the interactions weights and of their heterogeneity (see SM, Sec. IB), obtained considering the hypergraph aggregated from time 0 to time $t$. We note that the topology generated without the memory effects (EATHw) is very different: the average hyperdegree and projected degree are much higher, the distributions span a much broader interval, and the (hyper)degree reach in particular very large values; some nodes participate in a very large number of groups, having a large neighborhood, and only few groups and links are recurring (see Fig. 5 and SM, Sec. IB). The empirical higher-order structure is thus reproduced by the combination of the memory mechanisms and the heterogeneity in node activity, as also noted in [12]. Their combined effect also allows to reproduce modularity values of the pairwise projection of the aggregated hypergraph close to the ones observed in the data, while the EATHw model leads always to very small modularity values (see SM, Sec. IB).
@@
Figure 7: Generated nodes and hyperedges heterogeneity. Panels a and b show respectively the distributions of the instantaneous activity, $P(a_{h}(i))$, and of the persistence activity, $P(a_{T}(i))$, extracted from the empirical dataset (which are provided as input to the generation) and measured in the generated hypergraphs. In panel c (resp. f) each point corresponds to a node (resp. hyperedge) and we show the scatterplot between the burstiness of their inter-event time distributions, $\Delta B_{\tau}$, and of their event duration distributions, $\Delta B_{T}$. The gray dashed lines are the reference $\Delta B=0$. Panels d, e (resp. g, h) show the distributions $P(\Delta B_{T})$ and $P(\Delta B_{\tau})$ for nodes (resp. hyperedges): we consider only nodes and hyperedges with at least 1010 events. Note that in panels g and h there are no results for the EATHw case, since no group activated enough times to define its burstiness parameter. In all panels we consider the SFHH dataset and we compare it with hypergraphs generated using the model with (EATH) and without memory (EATHw), with the parameters extracted from the empirical hypergraph.
@@
Figure 6 focuses on the distributions of the quantities characterizing the dynamics of nodes and groups, at the global level: the dynamics generated for groups and nodes is heterogeneous, bursty and correlated, as empirically observed (see also the SM, Sec. IC, for the burstiness $\Delta B$ of the generated distributions). We note that, at the node level, the distributions of the number of events in trains of events and the inter-event times are well replicated even with the memory-less model EATHw. Indeed, the interaction dynamics of nodes is largely governed by their activity dynamics, which is built in the first step of the model (without reference to memory mechanisms). However, the EATHw generation process produces temporal dynamics that do not reproduce the empirical distributions of event durations for nodes nor any of the distributions characterizing the dynamics of groups (see Fig. 6). In this case indeed, the active groups are selected at each time-step only according to the activity of the nodes, thus it is not possible to obtain long durations for group interactions nor temporal correlations. Memory mechanisms are thus responsible for the generation of heterogeneous group durations, and allow also to reactivate the same groups at short intervals, creating trains of events [38, 43].
@@
Figure 8: Generated nodes participation to different group sizes. We consider the empirical SFHH hypergraph (Data) and the hypergraphs generated by the model with (EATH) and without memory (EATHw), with parameters extracted from the empirical hypergraph. In each case, we rank the nodes at each order $m$ according to the time spent interacting at that order. We then show the Pearson’s correlation coefficient $\rho(m,m^{\prime})$ between the rankings at order $m$ and $m^{\prime}$ (a,c,e). We show the evolution of the fraction $n_{f}^{m}(t)$ of active nodes among the ones occupying the top fNfN position of the ranking at order $m$, for different $m$ (see legend), and $f=0.1$, focusing on the same time window (b,d,f). We also plot the total fraction $N_{t}/N$ of active nodes in the population. Panel g shows the distributions of the node participation ratios $P(y)$ (the red dashed line represents y=1/(M1)y=1/(M-1)).
@@
Finally, we highlight a limitation of the model. Not surprisingly given the creation rules described above, the surrogate temporal hypergraphs generated do not reproduce the dynamics of group aggregation and disaggregation (see SM, Sec. IE): indeed the model’s group selection mechanisms only consider processes in which a group can lose or gain a single node. However, the empirical dynamics is more complex, with merging and splitting of groups of different sizes, as described in [11]. To reproduce such dynamics, additional mechanisms should be introduced in the model, with rules making it possible for a group at time t1t-1 to lose or gain several nodes at once.
@@ IV.3 Higher-order SIR dynamics
Figure 9: Higher-order SIR process. Panels a-c and d-f show respectively the epidemic final size $R_{\infty}$ and the basic reproduction number $R_{0}$ as a function of the epidemiological parameters $(\lambda,\mu)$. The results are obtained by averaging over 400 numerical simulations, fixing $\nu=4$, and considering the spreading over the SFHH empirical hypergraph (Data) and over the corresponding temporal hypergraphs generated using the model with (EATH) and without memory (EATHw). $C$ indicates the Canberra distance [52] between the matrices obtained by simulating the process on the data and on the corresponding synthetic hypergraphs. In panels g,h we show the fraction of infected $I(t)$ and recovered $R(t)$ nodes as a function of time, when a random initial seed is infected at $t_{0}=0$, and averaging the curves over 200 realizations: panel g, $(\lambda,\mu)=(1.44\,10^{-2},5.66\,10^{-4})$ (see the blue star in panels a-f); panel h, $(\lambda,\mu)=(3.64\,10^{-2},3.22\,10^{-4})$ (see the purple circle in panels a-f).
The goal of the EATH model is to produce surrogate temporal hypergraphs that can be used to simulate dynamical processes on hypergraphs. We therefore consider here one such process, namely a prototypical spreading process including higher-order mechanisms, and compare its dynamics and outcome when simulated on either the empirical or the surrogate hypergraphs. Specifically, we consider the higher-order SIR (susceptible-infected-removed) epidemic process based on a non-linear infection rate [47], which assumes a higher probability of contagion in case of simultaneous exposure to multiple sources of infection [47, 16]. Each node of the hypergraph can be susceptible, $S$, infected, $I$, or recovered, $R$. The dynamics evolves in discrete time, with a time-step $\delta t$ equal to the resolution of the hypergraph $\mathscr{H}=\{(\mathcal{V}_{t},\mathcal{E}_{t})\}_{t=1}^{\mathcal{T}/\delta t}$. At each time-step $t$, a susceptible node $j\in\mathcal{V}_{t}$ that is part of a hyperedge $e\in\mathcal{E}_{t}$ with $i_{e,t}$ infected nodes, gets infected with probability $(1-e^{-\lambda i_{e,t}^{\nu}})$ within the group $e$, where $\lambda\in[0,1]$ and $\nu\geq 1$. Hence, overall a susceptible node $j$ at time $t$ is infected with probability:
@@
Figure 9 shows the results of the simulations performed on both empirical and surrogate hypergraphs for the SFHH dataset (see SM, Sec. IF, for other datasets). The SIR dynamics on the EATH model replicates well the empirical one for all the parameters configurations, both in terms of the total impact of the epidemic $R_{\infty}$ and of the initial spreading $R_{0}$ (see Fig. 9a-f). Furthermore, the epidemic evolution on the EATH model is similar to the empirical one, replicating the timing and incidence levels of the epidemic peaks, despite a slight overestimation of the peak heights (see Fig. 9g,h). Simulations on the hypergraphs generated without memory mechanisms (EATHw) yield a much faster dynamics in the early stages, with a much higher first peak and overall a substantial overestimation of the impact of the spread.
@@ V EATH artificial and hybrid hypergraphs
Figure 10: Hybrid hypergraphs - mixing WS16 (system) and Utah_elem (population). We consider two hybrid hypergraphs (Hybrid_sub and Hybrid_uni) generated with the EATH model considering the properties of the datasets WS16 (conference context; $d_{1}$ - system properties) and Utah_elem (school context; $d_{2}$ - population properties). a-f: distributions of inter-event times, durations and numbers of events in a train of events, for nodes (a-c) and hyperedges (d-f); g,h: normalized hyperdegree distribution in the aggregated hypergraph, $P(D/\langle D\rangle)$, and normalized degree distribution in the projected aggregated graph, $P(D_{proj}/\langle D_{proj}\rangle)$; i-l: Pearson’s correlation coefficient $\rho(m,m^{\prime})$ between the node rankings at size $m$ and $m^{\prime}$, obtained considering the time spent by nodes interacting at each order; m-r: epidemic final-size $R_{\infty}$ and basic reproduction number $R_{0}$ as a function of the SIR epidemiological parameters $(\lambda,\mu)$, obtained in the same conditions of Fig. 9; s-t: scatterplots of the burstiness of inter-event time distributions $\Delta B_{\tau}$ and that of duration distributions $\Delta B_{T}$ for single nodes and hyperedges, considering only the nodes and hyperedges with at least 10 events.
The interest and use of the EATH model actually go beyond the generation of surrogate temporal hypergraphs. The model parameters and mechanisms governing the different generative phases can be tuned in order to generate hypergraphs with various properties. This opens the possibility to build temporal hypergraphs with arbitrary features, possibly even different from the empirical ones, and to test their impact, e.g. on dynamical processes. At the context/system level, we can for instance build a hypergraph: (i) with a different temporal activity dynamics, imposing an arbitrary modulation $\Lambda_{t}$; (ii) with groups of different sizes, by changing the distribution $\Psi(m)$; (iii) with different temporal and interaction scales, e.g. a longer time interval $\mathcal{T}$ or a higher connectivity $\langle E_{t}\rangle$. At the population level, we can moreover: (i) consider a larger population $N$; (ii) consider arbitrary distributions of the activities, $a_{h}(i)$ and $a_{T}(i)$; (iii) set different order propensities $\varphi_{i}(m)$. In the SM, Sec. III, we generate such cases and explore how these modifications affect the temporal, topological and dynamical features of the system.
@@
(i) Hybrid substitution (Hybrid_sub). The hypergraph is generated with the system parameters of $d_{1}$ and the population parameters of $d_{2}$. This case models a population of a specific environment (e.g. students of a school), which preserves its features but interacts in a different context (e.g. a conference), with different temporal patterns, activity levels and sizes of interactions. The connection between the system and population properties requires modifying the order propensity $\varphi_{i}(m)$ according to the maximum interactions sizes: if $M_{2}\geq M_{1}$, $\varphi_{i}(m)$ is kept as in $d_{2}$; if $M_{2}<M_{1}$ we first fix $\varphi_{i}(m)=\varphi_{i}(M_{2})\forall m\in[M_{2}+1,M_{1}]$, so that nodes preserve their propensity for interactions at large sizes and we then normalize the resulting $\varphi_{i}(m)$ on the interval $2,M_{1}]$. We keep the level of interactions $\langle E_{t}\rangle=\langle E_{t,1}\rangle$, even if $N_{2}\neq N_{1}$, to focus on the pure effect of a population that interacts in a different environment, possibly characterized by a different interaction level;
@@
In Fig. 10 we show the properties of hybrid hypergraphs generated by mixing a conference (WS16, $d_{1}$) and a school dataset (Utah_elem, $d_{2}$). The generated hypergraphs differ from the empirical ones, but they still present realistic statistical properties for all the measures considered: a heterogeneous and bursty dynamics with temporal correlations both at the node and group level (Fig. 10 a-f), a realistic shape of the hyperdegree and degree distributions in the aggregated hypergraph and in its pairwise projection (Fig. 10 g,h), heterogeneity in the behaviours of the nodes both in the propensity to interact at different orders (Fig. 10 i-l) and temporally (Fig. 10 s,t). The outcome of a SIR dynamics presents as well a realistic pattern even if some differences are observed with respect to the original dataset $d_{1}$ (Fig. 10 m-r): in Hybrid_sub the epidemic is less effective, since the average number of interactions per node is lower; in Hybrid_uni the dynamics is more similar, but still differs due to the presence of $d_{2}$ nodes with different behaviours. In the SM, Sec. II, we also consider other hybrid hypergraphs.
@@ VI Discussion
The EATH model can generate surrogate temporal hypergraphs emulating empirical systems and their different levels of heterogeneity. It captures and reproduces a wide set of empirical temporal and topological properties. Moreover, the EATH model is able to replicate the higher-order properties of the system that shape higher-order infection dynamics, generating thus surrogate temporal hypergraphs useful to investigate contagion processes. We note that we have here leveraged the fact that the model parameters for surrogates generation can be easily extracted from the empirical data to be emulated. It would in principle be possible to retrieve them using methods based on maximum likelihood. This represents a promising direction for future research.
@@ VII.1 Datasets and pre-processing
We consider several publicly available datasets of face-to-face interactions collected by different collaborations. The datasets consist in pairwise interactions collected with a resolution of 20 seconds through RFID proximity sensors. We consider six datasets collected by the SocioPatterns collaboration [49, 50, 51] describing interactions: in a hospital (LH10 [53], with $N=75$ and $\mathcal{T}=96$ hours), in an elementary school (LyonSchool [54], with $N=242$ and $\mathcal{T}=2$ days), in a high-school (Thiers13 [55], with $N=327$ and $\mathcal{T}=5$ days), in three conferences (SFHH [50], with $N=403$ and $\mathcal{T}=2$ days; ICCSS17 [56], with $N=274$ and $\mathcal{T}=4$ days; WS16 [56], with $N=138$ and $\mathcal{T}=2$ days); and two datasets collected by the Contacts among Utah’s School-age Population (CUSP) project [57], describing interactions in an elementary school (Utah_elem [57], with $N=339$ and $\mathcal{T}=2$ days) and in a middle school (Utah_mid [57], with $N=591$ and $\mathcal{T}=2$ days). These datasets describe social interactions in different environments, mediated by different mechanisms and activities.